Let fn(x) = cosnx/√ n , x belong to R, is not pointwise convergent. True or false with full explanation

1
Expert's answer
2022-03-17T07:02:51-0400

False.

Note that "\left| \frac{\cos nx}{\sqrt{n}} \right|\leqslant \frac{1}{\sqrt{n}}\rightarrow 0,n\rightarrow \infty"

Thus

"\underset{n\rightarrow \infty}{\lim}f_n\left( x \right) =\underset{n\rightarrow \infty}{\lim}\frac{\cos nx}{\sqrt{n}}=0"

which means "f_n" is pointwise convergent.